Should a nation tend to export or import goods for which it has a comparative advantage? Explain.

250 words.
No Plagiarism please.
Answer #1

The opportunity costs can be defined as the lost units of output of other goods for producing an additional unit of output of a good. A country has comparative advantage in the production of that good in which it has lower opportunity cost.

With the comparative advantage each country produces maximum goods by using its available resources efficiently. Hence total goods produced due to comparative advantage will be greater than the situation when every country produces every goods by using its resources inefficiently.

Hence due to comparative advantage extra output are produced and these extra goods go to both country due to trade. Hence comparative advantage leads to welfare increase of both country. The main benefit of free trade is that it provides more choices to consumers.

Let's take an example

Country A can make either 10 units of cloth or 1 units of wine in 10 hours.

Country B can make either 12 units of cloth or 6 units of wine in 10 hours.

Opportunity cost of 1 cloth for Country A is= 1/10

=0.1 wine

Opportunity cost of 1 wine for Country A is= 10/1

=10 clothe

Opportunity cost of 1 cloth for Country B is= 6/12

=0.5 wine

Opportunity cost of 1 wine for Country B is= 12/6

=2 clothes

Since opportunity cost of 1 units of cloth in Country A is 0.1 units of wine which is less than opportunity cost 1 units of cloth in Country B which is 0.5 units of wine.

Similarly opportunity cost of 1 units of wine in Country A is 10 units of clothes which is greater than opportunity cost 1 units of wine in Country B which is 2 units of clothes.

Hence Country B has a comparative advantage in wine production, Country A has a comparative in cloth production.